A 50.0 mg sample of an unknown radioactive substance was placed in storage and its mass measured periodically. After 70.9 days70.9 days, the amount of radioactive substance had decreased to 3.13 mg.

What is the half-life of the unknown radioactive substance?

half-life:

days
Answer #1

as w eknow from disintegration formula

ln ( N/No) = - Lt

L = ln ( 3.13/ 50) / 70.9

L = 0.03908 /day

expression for half life

t(1/2) = ln (2)/ L

t(1/2) = 17.735 days

======